OpenStudy (anonymous):

Write the complex number in the form a + bi. sqrt6(cos 315° + i sin 315°)

OpenStudy (anonymous):

\[\sqrt{6}(\cos 315^{o}+i \sin315^{o})\] \[\sqrt{6}(\frac{ \sqrt{2} }{ 2 }-\frac{ \sqrt{2} }{ 2 }i)\] \[\frac{ \sqrt{12}}{ 2 }-\frac{ \sqrt{12} }{ 2 }i\] \[\frac{ 2\sqrt{3} }{ 2 }-\frac{2\sqrt{3} }{2 }i\] \[\sqrt{3}-\sqrt{3}i\]